# references.bib

### v2

@@ 4.1 Policy Evaluation: RISE-2
To evaluate the performance of policies in both in-domain and generalization settings, we design 4 tasks in Fig. 3 and collect 50 demonstrations for each task using teleoperation [airexo_v1]. We compare RISE-2 against several representative policies based on 2D images or 3D point clouds, including ACT [act], Diffusion Policy [dp], CAGE [cage], RISE [rise], and $\pi_{0}$ [pi0]. Additionally, we include a RISE-2 variant with the dense encoder implemented with an ImageNet-pretrained [imagenet] ResNet-18 [resnet], denoted as RISE-2 (ResNet-18). Please refer to Appendix C.1 for details.
@@
As shown in Fig. 3, RISE-2 significantly outperforms previous state-of-the-art imitation policies in average success rate across 4 tasks, demonstrating its effectiveness in handling diverse manipulation challenges. On tasks requiring precise motion execution, such as Lift Plate and Open Lid, RISE-2 achieves superior accuracy in predicting fine-grained actions, consistently surpassing all baselines. The performance gain over RISE [rise] highlights the benefits of using separate 2D and 3D encoders with spatial feature alignment. Further improvements from RISE-2 (ResNet-18) to RISE-2 (DINOv2) underscore the value of leveraging diverse knowledge from visual foundation models [dino] in robotic manipulation.
@@
ACT [act] 32.5% 0.0% 2.5% 0.0%
Diffusion Policy [dp] 40.0% 12.5% 5.0% 2.5%
CAGE [cage] 65.0% 45.0% 42.5% 10.0%
RISE [rise] 72.5% 42.5% 12.5% 40.0%
$\pi_{0}$ [pi0] 85.0% 77.5% 82.5% 82.5%
@@ 4.2 System Evaluation: AirExo-2
Previous experiments have demonstrated that RISE-2 is a generalizable policy, making it well-suited for learning from in-the-wild demonstrations collected and processed by AirExo-2. In this evaluation, we utilize AirExo-2 to collect an equal number of in-the-wild demonstrations and convert them into pseudo-robot demonstrations. These demonstrations are then used to train the RISE-2 policy, which is deployed on a real robot without further fine-tuning. As shown in Fig. 4, the policy achieves satisfactory success rates on all tasks, with only a slight performance drop compared to the policy trained on teleoperated demonstrations. This highlights the ability of the AirExo-2 system to collect and adapt high-quality demonstrations suitable for training downstream imitation policies. Please refer to Appendix C.4 for additional results of the RISE policy [rise].
@@
We use the Collect Toys task to highlight the importance of the demonstration adaptor for learning from in-the-wild data. As shown in Tab. 2, removing visual adaptors significantly reduces performance, revealing a substantial visual gap between in-the-wild and robot demonstrations. These results also validate the quality of AirExo-2 pseudo-robot demonstrations for training high-performing policies. Notably, RISE-2 without adaptors still performs comparably to RISE [rise] with adaptors, demonstrating its strong generalization across embodiments.
@@ 4.3 System Analysis: AirExo-2
Method Completion Time (s) $\downarrow$ Avg. Rank $\downarrow$ Rating $\uparrow$
@@
We compare three demonstration collection methods: end-effector pose teleoperation using a haptic device [rh20t], joint-space teleoperation via AirExo [airexo_v1], and in-the-wild collection with AirExo-2. Twenty participants (14 men, 6 women, aged 21–35) of varying expertise levels perform the Collect Toys task on all three platforms in randomized order, with each trial consisting of a 3-minute familiarization and a timed demonstration collection. After completing all trials, participants fill out a questionnaire to provide feedback on three collection methods. From the results in Tab. 3, we observe that both experienced and novice users find AirExo-2 easier to learn and operate than both types of teleoperation, leading to faster onboarding and smoother operation. Statistical analysis (Welch’s $t$-test, $p=8.32\times 10^{-10}$) further confirms its superior efficiency in task completion time. The user-friendliness and efficiency make AirExo-2 well-suited for diverse users and consistent, high-quality data collection at scale in real-world settings.
Device Error (mm) $\downarrow$
Avg ±${}_{\pm\ \text{Std}}$ Max
UMI [umi] $8.855_{\pm 3.228}$ 20.002
@@
Table 4: Action Accuracy of Data Collection Systems. AirExo-2 achieves higher action accuracy than handheld devices like UMI [umi].
@@
Accurate actions are essential for learning reliable policies from in-the-wild demonstrations. Unlike handheld devices such as UMI [umi], which use visual SLAM for trajectory estimation, AirExo-2 leverages its mechanical structure and forward kinematics for more precise motion tracking. As shown in Tab. 4, it achieves significantly lower translation errors across 3 tracks (please refer to Appendix C.2 for details about accuracy evaluations), confirming its reliability for high-fidelity motion capture.
@@ 6 Limitations and Future Works
While we utilize the proposed demonstration adaptor to visually transform in-the-wild demonstrations collected by AirExo-2 into pseudo-robot demonstrations, these transformed demonstrations are primarily useful for generalizable policies. To enhance the applicability of the pseudo-robot demonstrations to a broader range of policies, future work could explore the integration of demonstration augmentation methods, such as novel view synthesis [rovi_aug, zeronvs, vista, gcd], into the adaptation process to improve the diversity of the demonstrations, making them more versatile for various policy learning.
As demonstrated by several works [umi, inhand_than_global, giving_robots_a_hand, demoat], the in-hand image is a semi-unified observation modality across different embodiments. Our case study also reveals that combining in-hand observations can strengthen the performance of various 2D policies. However, the current AirExo-2 system does not include in-hand cameras. Although we have designed connectors to integrate them with the exoskeleton, calibrating the in-hand cameras with the exoskeleton remains challenging, making it difficult to adapt the in-hand images into the robot domain. Future work could explore effective methods for adapting in-hand images collected by AirExo-2 to the robot domain, or investigate strategies for leveraging the semi-unified in-hand observations in policy design.
The current AirExo-2 system only supports parallel grippers as end-effectors, limiting its applicability in more dexterous tasks. Future work could integrate the AirExo-2 system with dexterous hands [eyesight_hand, leap_hand] and their corresponding exoskeletons [ben2025homie, other_hand], enabling more complex manipulation capabilities and expanding the range of tasks the system can effectively perform. Moreover, adding additional sensors to the AirExo-2 system for recording other observation modalities — such as force/torque [foar, forcemimic], tactile [vitamin], and sound [maniwav] — can also be explored in the future.
[…]
@@ A.1 Hardware Design
High-quality in-the-wild data requires precise collection for effective pseudo-robot transformation. However, AirExo [airexo_v1] has several key hardware limitations:
@@
Following [umi, forcemimic], the gripper of AirExo-2 incorporates a linear guide, with the fingers mounted on a sliding block. This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly under clamping forces without any stalling.
@@ A.2 Calibration
For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining $\{\tilde{q}_{\text{calib}}^{\text{left}},\tilde{q}_{\text{calib}}^{\text{right}}\}$. The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base $\mathbf{T}^{\text{base}}_{\text{marker}}$ and performing optical calibration using OpenCV [opencv], obtaining $\mathbf{T}^{\text{camera}}_{\text{marker}}$. Thus, the transformation can calculated as
@@
In the second stage, inspired by prior works [easyhec, easyhec++, sam-rl], we use differentiable rendering [differentiable_rendering] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [redner]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [sam2]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy.
@@
where parameters, except for the base-to-camera rotation, are represented as deltas relative to the initial calibration results. The base-to-camera rotation is expressed in a 6D format [rot6d]. Thus, the final calibration results can be calculated as
@@
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as $\{I_{i},d_{i},q_{i}^{\text{left}},q_{i}^{\text{right}}\}_{i=1}^{N_{c}}$, where $N_{c}$ represents the total number of training samples for calibration, and $I_{i}$ and $d_{i}$ are the RGB and depth images of the $i$-th sample, respectively. Subsequently, we utilize SAM-2 [sam2] to annotate the AirExo-2 mask $M^{a}_{i}$ and the depth mask $M^{d}_{i}\subseteq M^{a}_{i}$ for each sample $i$, as shown in Fig. 8. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
@@
The differentiable rendering engine Redner [redner] is employed to render the AirExo-2 mask $\hat{M}_{i}^{a}$ and AirExo-2 depth $\hat{d}_{i}$ using the calibration results and joint information:
@@
where $\beta$ represents the weighting coefficient, and $\circ$ denotes the mask-apply operation. In practice, we set $\beta=5$, use $N_{c}=40$ samples for optimization, and employ the Adam optimizer [adam] with a learning rate of $10^{-4}$ for 1000 iterations to fine-tune the calibration results.
@@ A.3 Visual Adaptors
In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [sam2]. After a few annotations on approximately 50 scenes, we can fine-tune SAM-2 on human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptation process, making it nearly fully automated.
@@
We train a ControlNet [controlnet] based on the Stable Diffusion 1.5 [stable_diffusion] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions.
@@
For training, we use a batch size of 88 and a learning rate of $10^{-5}$, while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [ddpm] with a guidance scale of 9.0 [cfg]. The prompt for generating robot images for our robot platform is displayed as follows.
@@ B.1 Data Processing
The color image is resized to 448$\times$252 for DINOv2 backbone [dino] and 640$\times$360 for ResNet-18 backbone [resnet]. The depth image is resized to 640$\times$360 before creating the point cloud. The camera intrinsics are adjusted accordingly. Both the point clouds and actions are projected to the camera coordinate system following [rise, idp3].
@@ B.2 Hyperparameters
The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [minkowski]. The dense encoder adopts DINOv2-base [dino] as the 2D backbone with the output channel of 128. In the spatial aligner, we use $M=3$ for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. 6. The transformer in action generator contains 4 blocks, in which we set $d_{\mathrm{model}}=512$ and $d_{\mathrm{ff}}=2048$ following [rise]. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [dp] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20.
@@
Table 6: Sparse Convolutional Network Parameters of the RISE-2 Policy. Both the sparse encoder and the spatial aligner utilize MinkResNet [minkowski] for point cloud encoding. $k,c,d,s$ stand for the kernel size, output channel number, dilation, and stride in the convolutional layers, respectively.

### v3

@@ 4.1 Policy Evaluation: RISE-2
To evaluate the performance of policies in both in-domain and generalization settings, we design 4 tasks in Fig. 3 and collect 50 demonstrations for each task using teleoperation [29]. We compare RISE-2 against several representative policies based on 2D images or 3D point clouds, including ACT [123], Diffusion Policy [18], CAGE [111], RISE [102], and $\pi_{0}$ [7]. Additionally, we include a RISE-2 variant with the dense encoder implemented with an ImageNet-pretrained [22] ResNet-18 [33], denoted as RISE-2 (ResNet-18). Please refer to Appendix C.1 for details.
@@
As shown in Fig. 3, RISE-2 significantly outperforms previous state-of-the-art imitation policies in average success rate across 4 tasks, demonstrating its effectiveness in handling diverse manipulation challenges. On tasks requiring precise motion execution, such as Lift Plate and Open Lid, RISE-2 achieves superior accuracy in predicting fine-grained actions, consistently surpassing all baselines. The performance gain over RISE [102] highlights the benefits of using separate 2D and 3D encoders with spatial feature alignment. Further improvements from RISE-2 (ResNet-18) to RISE-2 (DINOv2) underscore the value of leveraging diverse knowledge from visual foundation models [75] in robotic manipulation.
@@
ACT [123] 32.5% 0.0% 2.5% 0.0%
Diffusion Policy [18] 40.0% 12.5% 5.0% 2.5%
CAGE [111] 65.0% 45.0% 42.5% 10.0%
RISE [102] 72.5% 42.5% 12.5% 40.0%
$\pi_{0}$ [7] 85.0% 77.5% 82.5% 82.5%
@@ 4.2 System Evaluation: AirExo-2
Previous experiments have demonstrated that RISE-2 is a generalizable policy, making it well-suited for learning from in-the-wild demonstrations collected and processed by AirExo-2. In this evaluation, we utilize AirExo-2 to collect an equal number of in-the-wild demonstrations and convert them into pseudo-robot demonstrations. These demonstrations are then used to train the RISE-2 policy, which is deployed on a real robot without further fine-tuning. As shown in Fig. 4, the policy achieves satisfactory success rates on all tasks, with only a slight performance drop compared to the policy trained on teleoperated demonstrations. This highlights the ability of the AirExo-2 system to collect and adapt high-quality demonstrations suitable for training downstream imitation policies. Please refer to Appendix C.4 for additional results of the RISE policy [102].
@@
We use the Collect Toys task to highlight the importance of the demonstration adaptor for learning from in-the-wild data. As shown in Tab. 2, removing visual adaptors significantly reduces performance, revealing a substantial visual gap between in-the-wild and robot demonstrations. These results also validate the quality of AirExo-2 pseudo-robot demonstrations for training high-performing policies. Notably, RISE-2 without adaptors still performs comparably to RISE [102] with adaptors, demonstrating its strong generalization across embodiments.
@@ 4.3 System Analysis: AirExo-2
Method
Completion
Time (s) $\downarrow$
Avg. Rank $\downarrow$ Rating $\uparrow$
@@
We compare three demonstration collection methods: end-effector pose teleoperation using a haptic device [27], joint-space teleoperation via AirExo [29], and in-the-wild collection with AirExo-2. Twenty participants (14 men, 6 women, aged 21–35) of varying expertise levels perform the Collect Toys task on all three platforms in randomized order, with each trial consisting of a 3-minute familiarization and a timed demonstration collection. After completing all trials, participants fill out a questionnaire to provide feedback on three collection methods. From the results in Tab. 3, we observe that both experienced and novice users find AirExo-2 easier to learn and operate than both types of teleoperation, leading to faster onboarding and smoother operation. Statistical analysis (Welch’s $t$-test, $p=8.32\times 10^{-10}$) further confirms its superior efficiency in task completion time. The user-friendliness and efficiency make AirExo-2 well-suited for diverse users and consistent, high-quality data collection at scale in real-world settings.
Device Error ($m$) $\downarrow$
Avg $_{\pm\ \text{Std}}$ Max
UMI [19] $8.855_{\pm 3.228}$ 20.002
@@
Table 4: Action Accuracy of Data Collection Systems. AirExo-2 achieves higher action accuracy than handheld devices like UMI [19].
@@
Accurate actions are essential for learning reliable policies from in-the-wild demonstrations. Unlike handheld devices such as UMI [19], which use visual SLAM for trajectory estimation, AirExo-2 leverages its mechanical structure and forward kinematics for more precise motion tracking. As shown in Tab. 4, it achieves significantly lower translation errors across 3 tracks (please refer to Appendix C.2 for details about accuracy evaluations), confirming its reliability for high-fidelity motion capture.
@@ 6 Limitations and Future Works
While we utilize the proposed demonstration adaptor to visually transform in-the-wild demonstrations collected by AirExo-2 into pseudo-robot demonstrations, these transformed demonstrations are primarily useful for generalizable policies. To enhance the applicability of the pseudo-robot demonstrations to a broader range of policies, future work could explore the integration of demonstration augmentation methods, such as novel view synthesis [14, 85, 94, 95], into the adaptation process to improve the diversity of the demonstrations, making them more versatile for various policy learning.
As demonstrated by several works [19, 41, 54, 115], the in-hand image is a semi-unified observation modality across different embodiments. Our case study in Appendix C.3 also reveals that combining in-hand observations can strengthen the performance of various 2D policies. However, the current AirExo-2 system does not include in-hand cameras. Although we have designed connectors to integrate them with the exoskeleton, calibrating the in-hand cameras with the exoskeleton remains challenging, making it difficult to adapt the in-hand images into the robot domain. Future work could explore effective methods for adapting in-hand images collected by AirExo-2 to the robot domain, or investigate strategies for leveraging the semi-unified in-hand observations in policy design.
The current AirExo-2 system only supports parallel grippers as end-effectors, limiting its applicability in more dexterous tasks. Future work could integrate the AirExo-2 system with dexterous hands [84, 89] and their corresponding exoskeletons [28, 3], enabling more complex manipulation capabilities and expanding the range of tasks the system can effectively perform. Moreover, adding additional sensors to the AirExo-2 system for recording other observation modalities — such as force/torque [34, 63], tactile [61], and sound [64] — can also be explored in the future.
Acknowledgments
We would like to thank Peishen Yan and Chunyu Xue from Shanghai Jiao Tong University for their valuable suggestions on the figures in this paper. We would like to thank Teng Hu from Shanghai Jiao Tong University for insightful discussions about inpainting strategies to mitigate the visual gap.
This work was supported by the National Key Research and Development Project of China (No. 2022ZD0160102), the National Key Research and Development Project of China (No. 2021ZD0110704), Shanghai Artificial Intelligence Laboratory, XPLORER PRIZE grants.
@@ A.1 Hardware Design
High-quality in-the-wild data requires precise collection for effective pseudo-robot transformation. However, AirExo [29] has several key hardware limitations:
@@
Following [19, 63], the gripper of AirExo-2 incorporates a linear guide, with the fingers mounted on a sliding block. This design allows for smoother opening and closing of the gripper, ensuring it operates seamlessly under clamping forces without any stalling.
@@ A.2 Calibration
For initial calibration, the former calibration can be achieved by manually adjusting the joints to approximate the zero position using specialized 3D-printed tools and reading the encoder values, obtaining $\{\tilde{q}_{\text{calib}}^{\text{left}},\tilde{q}_{\text{calib}}^{\text{right}}\}$. The latter calibration can be done by attaching an ArUco calibration marker board with a known position on the base $\mathbf{T}^{\text{base}}_{\text{marker}}$ and performing optical calibration using OpenCV [8], obtaining $\mathbf{T}^{\text{camera}}_{\text{marker}}$. Thus, the transformation can calculated as
@@
In the second stage, inspired by prior works [15, 39, 66], we use differentiable rendering [51] to refine the initial calibration. Training samples are obtained from a single human play trajectory with AirExo-2. Using the joint states and calibration parameters, we render the system mask and depth via a differentiable rendering engine [58]. Calibration parameters are optimized by minimizing discrepancies between the rendered and annotated system masks, as well as between the rendered and observed depths. Pseudo-ground-truth masks are manually annotated with SAM-2 [82]. This iterative refinement compensates for errors that accumulate across joints, ultimately improving the overall accuracy.
@@
where parameters, except for the base-to-camera rotation, are represented as deltas relative to the initial calibration results. The base-to-camera rotation is expressed in a 6D format [126]. Thus, the final calibration results can be calculated as
@@
For the optimization process, we first record a single in-the-wild play trajectory, in which the human operator uses the AirExo-2 to adopt various poses. During trajectory recording, ensure that all parts of the AirExo-2 remain above the human hands and arms from the camera’s perspective. After data collection, we sample approximately 40 image-joint pairs from the trajectory, denoted as $\{I_{i},d_{i},q_{i}^{\text{left}},q_{i}^{\text{right}}\}_{i=1}^{N_{c}}$, where $N_{c}$ represents the total number of training samples for calibration, and $I_{i}$ and $d_{i}$ are the RGB and depth images of the $i$-th sample, respectively. Subsequently, we utilize SAM-2 [82] to annotate the AirExo-2 mask $M^{a}_{i}$ and the depth mask $M^{d}_{i}\subseteq M^{a}_{i}$ for each sample $i$, as shown in Fig. 8. The first mask provides supervision for the rendered AirExo-2 mask, while the second mask is used to select the valid depth information that serves as the supervision signal.
@@
The differentiable rendering engine Redner [58] is employed to render the AirExo-2 mask $\hat{M}_{i}^{a}$ and AirExo-2 depth $\hat{d}_{i}$ using the calibration results and joint information:
@@
where $\beta$ represents the weighting coefficient, and $\circ$ denotes the mask-apply operation. In practice, we set $\beta=5$, use $N_{c}=40$ samples for optimization, and employ the Adam optimizer [56] with a learning rate of $10^{-4}$ for 1000 iterations to fine-tune the calibration results.
@@ A.3 Visual Adaptors
In the image adaptor, we initially annotate the hand mask (and, if visible, the head mask) manually using SAM-2 [82]. After a few annotations on approximately 50 scenes, we can fine-tune SAM-2 on human-annotated samples, enabling automated labeling. This significantly reduces human effort and streamlines the demonstration adaptation process, making it nearly fully automated.
@@
We train a ControlNet [120] based on the Stable Diffusion 1.5 [83] model to generate photo-realistic robot images from rendered robot images. To collect training samples, we use teleoperation to gather a small amount of play data, where the robot is teleoperated to move randomly within an empty workspace while recording RGB-D images and corresponding joint states. This ensures a diverse dataset of robot arm configurations, free from occlusions or distractions.
@@
For training, we use a batch size of 88 and a learning rate of $10^{-5}$, while keeping other hyperparameters at their default settings. We use 50 DDPM sampling steps [37] with a guidance scale of 9.0 [38]. The prompt for generating robot images for our robot platform is displayed as follows.
@@
A.4 Discussions
Hardware Improvements over AirExo [29].
We improved both the repairability and accuracy of the system to enable scalable, high-quality data collection. In terms of repairability, relocating the encoders outside the joint and adopting a gear-driven design reduced encoder replacement time from 15 minutes to just 2 minutes. In terms of accuracy, the system’s action recording precision was significantly enhanced, as presented in Tab. 4.
Computation Overhead of Adaptation.
Rendering a 1-minute scene of approximately 600 frames requires about 40 seconds on a single GPU, with the majority of the time consumed by the ControlNet-tuned Stable Diffusion [83, 120] model. This process can be further accelerated by: (1) rendering at a lower resolution, and (2) leveraging multiple GPUs in parallel. Notably, the primary bottleneck in data collection lies in human effort rather than computation. Given that computational resources can be scaled more readily, we argue that prioritizing the reduction of human involvement — even at the cost of slightly increased computational usage — can significantly improve the overall efficiency and scalability of the data collection pipeline.
Cross-Robot Transfer and Data Reusability.
Our adaptation pipeline can be easily extended to other robots by deriving joint values from end-effector poses via inverse kinematics, followed by rendering and inpainting the corresponding robot. As shown in Tab. 2, even without visual adaptation, policies trained solely on in-the-wild demonstrations can sometimes complete tasks successfully. This suggests that raw in-the-wild data itself can already be used to train policies and deployed on different robots. By incorporating robot-specific adaptations, we can generate high-quality data for different robots and boost policy performance by training on these adapted demonstrations. This approach can greatly enhance data reusability across platforms and embodiments.
Strict 1:1 Joint Mapping Design of AirExo-2.
In our experiments, the Flexiv arms are comparable in size to the commonly used robotic arms like Franka, KUKA, and xArms. As shown in Tab. 3, the user study indicates that operators find our system intuitive and easy to use, suggesting that exact 1:1 scaling is not a prerequisite for effective data collection. Moreover, as most humanoid robots are designed to match human dimensions, our system is naturally compatible. For manipulators of different sizes, human-sized exoskeletons can still be employed, with data transfer facilitated through the adaptation process described in the previous discussion.
@@ B.1 Data Processing
[…]
@@ B.2 Hyperparameters
The sparse encoder adopts a ResNet-like architecture built upon MinkowskiEngine [20]. The dense encoder adopts DINOv2-base [75] as the 2D backbone with the output channel of 128. In the spatial aligner, we use $M=3$ for feature alignment. The aligned features are fused by shared MLPs with the size of (256, 256, 256), and then fed into another sparse network. The two sparse networks are detailed in Tab. 6. The transformer in action generator contains 4 blocks, in which we set $d_{\mathrm{model}}=512$ and $d_{\mathrm{ff}}=2048$ following [102]. The channel number of the readout token is 512. The diffusion head adopts a CNN implementation [18] with 100 denoising iterations in training and 20 iterations in inference. The output action horizon used in experiments is 20.
@@
Table 6: Sparse Convolutional Network Parameters of the RISE-2 Policy. Both the sparse encoder and the spatial aligner utilize MinkResNet [20] for point cloud encoding. $k,c,d,s$ stand for the kernel size, output channel number, dilation, and stride in the convolutional layers, respectively.
@@
Inference.
We deploy RISE-2 using an RTX 2060 SUPER GPU and achieve a 5Hz inference speed, demonstrating efficient performance compared to larger vision-language-action policies [7]. We additionally provide scripts that enable policy inference to be executed on a remote server.